OpenStudy (anonymous):

Math Analysis: Use fundamental identities to simplify the expression. 14. sin^2 x + sinx - 1 + cos^2 x 15. cotx sinx - tanxcosx

OpenStudy (apoorvk):

\[sin^2x + cos^2x = 1\] use this identity in the first one. \[cotx = \frac {cosx}{sinx}~~and ~~tanx=\frac {sinx}{cosx}\] use this in the second one.

OpenStudy (btaylor):

14. sin^2 x + cos^2 x = 1, so 1 - 1 + sin x = sin x 15. cot x = (cos x)/(sin x), so cotx sinx = cos x. tan x = (sin x)/(cos x), so tanx cosx= sin x. So the end result is cos x - sin x

OpenStudy (anonymous):

#15. so far i got cos x - sin x what do i do next?

OpenStudy (apoorvk):

that's it. you let it remain that way for #15. That's as simplified as it can get.